# Trees of Graphs as Boundaries of Hyperbolic Groups

### v1

@@ 3.1. The Bowditch JSJ splitting of a 1-ended hyperbolic group
Naturality of the homeomorphism $\partial G\cong\lim\Theta(G)$ in the above statement means that the inclusions $\partial G(t)\subset\partial G$ and $\partial G(u)\subset\partial G$ induce a map $\lim\Theta(G)\to\partial G$ which is a homeomorphism.
@@ 3.2. The reduced Bowditch JSJ splitting and its rigid cluster factors
Let $G$ be a 1-ended hyperbolic group which is not cocompact Fuchsian and let ${T}_{G}$ be the Bowditch JSJ tree of $G$. Let $\sim$ be the smallest equivalence relation on the vertex set of ${T}_{G}$ for which $u\sim v$ whenever $u$ and $v$ are rigid vertices at distance $2$. Equivalence classes of $\sim$ are either white vertex singletons, flexible vertex singletons or rigid classes (i.e. sets of rigid vertices). The convex hulls of the rigid classes are the rigid clusters of ${T}_{G}$, and they have a form of subtrees of $T_{G}$ (possibly reduced to singletons).
@@
We now describe closer the nature of the just defined reduced Bowditch JSJ splitting, and in particular of its factors. Observe that the bipartite structure of the tree ${T}_{G}$ descends to ${T}^{r}_{G}$, so that we can speak of black and white vertices of the latter. The white vertices of ${T}^{r}_{G}$ are in the bijective correspondence with those white vertices of ${T}_{G}$ which are adjacent to at least one flexible vertex. Their stabilizers (for the induced action of $G$ on ${T}^{r}_{G}$) coincide with the corresponding stabilizers for the $G$-action on ${T}_{G}$, and so they are maximal 2-ended subgroups of $G$. There are two kinds of black vertices in ${T}^{r}_{G}$. The first kind black vertices of ${T}^{r}_{G}$ are in the bijective correspondence with the flexible vertices of ${T}_{G}$, and so we still call them flexible. Their stabilizers under the action of $G$ on ${T}^{r}_{G}$ do not change, and we call them the flexible factors of the reduced Bowditch JSJ splitting of $G$. The second kind black vertices of ${T}^{r}_{G}$ correspond bijectively to the rigid clusters of ${T}_{G}$, and we call them the rigid cluster vertices of ${T}^{r}_{G}$. Their stabilizers, which coincide with the stabilizers of the corersponding rigid clusters in the $G$-action on ${T}_{G}$, are called the rigid cluster factors of the reduced Bowditch JSJ splitting of $G$.
Next observations follow directly from the above description.
@@
Each rigid cluster vertex $v$ of the reduced Bowditch tree ${T}^{r}_{G}$ is isolated, i.e. there is no rigid cluster vertex in ${T}^{r}_{G}$ at distance 2 from $v$. Equivalently, all vertices of ${T}^{r}_{G}$ lying at distance 2 from a rigid cluster vertex $v$ are flexible.
@@
Given a rigid cluster $S\subset{T}_{G}$, consider the action of the corresponding rigid cluster factor $G_{S}$ on $S$. The vertex stabilizers of this action are easily seen to coincide with the corresponding vertex stabilizers of the $G$-action on ${T}_{G}$. Thus, this action describes some splitting of the rigid cluster factor $G_{S}$, along its maximal 2-ended subgroups, whose factors are some appropriate rigid factors of the Bowditch JSJ splitting of $G$. This observation helps to understand the nature of rigid cluster factors, even though we will make no essential use of it.
Since the factors of any splitting of a hyperbolic group along its quasi-convex subgroups are also quasi-convex (see e.g. Proposition 1.2 in [1]), and since 2-ended subgroups are quasi-convex, we get the following.
@@
where each $\varphi_{e}:\partial G({w(e)})\to\partial G({b(e)})$ is the natural inclusion. The comments from the paragraph preceding the statement of Theorem 3.2 yield also the following result.
@@
Let $G$ be a 1-ended hyperbolic group which is not cocompact Fuchsian, and let $\Theta^{r}(G)$ be the associated tree system (related to the reduced Bowditch JSJ splitting of $G$). Then the Gromov boundary $\partial G$ is naturally homeomorphic to the limit of the tree system $\Theta^{r}(G)$, i.e.
@@ 4. General trees of graphs
In this section we describe trees of graphs - a class of topological spaces which is the main object of study in this paper. As we will see, these spaces are compact metrisable, and the most interesting subclass (that of trees of 2-connected graphs) consists of spaces that are connected, locally connected, cutpoint-free, and of topological dimension 1. Throughout all of this section we make use of the terminology and notation introduced in Section 2. We start with basic definitions.
@@
for a tree system of punctured graphs (where for each $t\in V^{b}_{T}$, $\Gamma_{t}$ is some finite essential graph).
@@ 4.1. Trees of graphs as inverse limits of graphs
We will now present an alternative description of trees of graphs, as inverse limits of some special inverse systems of graphs. (As we will see below, these inverse systems in fact coincide, up to isomorphism, with the systems ${\mathcal{S}}_{\Theta}$ as described in Section 2.6, associated to the corresponding tree systems $\Theta$ of punctured graphs.) This alternative approach to trees of graphs is related to (and generalizes) that given in Section 2 of [24], where only a special case of the so called reflection trees of graphs has been treated, using similar description.
Let $\Theta$ be a tree system of punctured graphs. For any graph $\Gamma_{t}$ whose punctured version $\Gamma_{t}^{\circ}$ appears as a constituent space in $\Theta$, denote by $D_{t}\subset|\Gamma_{t}|$ a countable dense subset of the realization $|\Gamma_{t}|$ containing all vertices of $\Gamma_{t}$. Note that such $D_{t}$ is unique up to a homeomorphism of $|\Gamma_{t}|$ identical on the vertices and preserving all the edges. Consider also the inverse system ${\mathcal{S}}_{\Theta}=[\{K_{F}^{\ast}\},\{f_{F_{1}F_{2}}\}]$ associated to $\Theta$. Observe that, viewing any black vertex $t\in V^{b}_{T}$ as a b-subtree of $T$, and denoting by $\{[\varphi_{[u,t]}(\Sigma_{u})]:u\in N_{t}\}$ the set of points of the quotient space $K_{t}^{\ast}$ obtained from $\Gamma_{t}^{\circ}$ by shrinking the peripheral sets $\varphi_{[u,t]}(\Sigma_{u}):u\in N_{t}$, there is a homeomorphism $h_{t}:K_{t}^{\ast}\to|\Gamma_{t}|$ such that:
@@
    $h_{t}$ respects types of the shrinked peripheral sets $[\varphi_{[u,t]}(\Sigma_{u})]$, i.e. if $\varphi_{[u,t]}(\Sigma_{u})$ is the boundary of a normal neighbourhood of some vertex of $\Gamma$, then $h_{t}([\varphi_{[u,t]}(\Sigma_{u})])$ is equal to this vertex, and if $\varphi_{[u,t]}(\Sigma_{u})$ is the boundary of a normal neighbourhood of some interior point of some edge $a$ of $\Gamma$, then $h_{t}([\varphi_{[u,t]}(\Sigma_{u})])$ is equal to some interior point of the same edge $a$.
@@
Let $X$ be the realization of a finite graph, and let $x\in X$. The blow-up of $X$ at $x$ is the space obtained from $X$ by deleting $x$, and by taking the completion of the resulting space which consists of adding as many points as the number of connected components into which $x$ splits its normal neighbourhood $B(x)$ in $X$. We denote this blow-up by $X^{\#}(x)$, and the set of points added at the completion by $P_{x}$ (we use the term blow-up divisor at $x$ for this set $P_{x}$). Similarly, if $J$ is a finite subset of $X$, the blow-up of $X$ at $J$, denoted $X^{\#}(J)$, is the space obtained from $X$ by performing blow-ups at all points of $J$ (in arbitrary order). Moreover, for any two finite subsets $J_{1}\subset J_{2}$ of $X$, denote by $\rho^{X}_{J_{2}J_{1}}:X^{\#}(J_{2})\to X^{\#}(J_{1})$ the blow-down map which maps the sets $P_{x}:x\in J_{2}\setminus J_{1}$ to the corresponding points $x$, and which is identical on the remaining part of $X^{\#}(J_{2})$.
Now, let $F$ be a finite b-subtree of the underlying tree $T$ of the tree system $\Theta$. For each $t\in V^{b}_{F}$ consider the subset $J_{t}^{F}\subset|\Gamma_{t}|$ given by $J_{t}^{F}:=\{h_{t}([\varphi_{[t,u]}(\Sigma_{u})]):u\in N^{F}_{t}\}$. Define $X_{F}$ as quotient of the disjoint union
@@
through the equivalence relation $\sim$ determined by the following gluing maps. Recall that, by condition (2) of Definition 4.1, for any white vertex $u$ of $F$ there are exactly two black vertices in $F$ adjacent to $u$. Denoting these adjacent black vertices by $t$ and $s$, consider the blow-up divisors $P_{h_{t}([\varphi_{[t,u]}(\Sigma_{u})])}\subset|\Gamma|^{\#}(J_{t}^{F})$ and $P_{h_{s}([\varphi_{[s,u]}(\Sigma_{u})])}\subset|\Gamma|^{\#}(J_{s}^{F})$. Note that there is a natural bijection

### v2

@@ 3.1. The Bowditch JSJ splitting of a 1-ended hyperbolic group
The homeomorphism $\partial G\cong\lim\Theta(G)$ in the above statement is natural in the sense that it is induced by the inclusions $\partial G(t)\subset\partial G$ and $\partial G(u)\subset\partial G$, where $t\in V_{{\mathcal{T}}_{G}}^{b}$, $u\in V_{{T}_{G}}^{w}$.
Proof.
We sketch here an explanation of how Theorem 3.2 follows from the result of Alexandre Martin mentioned right before the statement of this theorem.
In Section 2.6.2 of [CM17] Alexandre Martin and Christopher Cashen consider an object $X$ called a proper geodesic hyperbolic tree of spaces, and the associated object $\partial X$ (which may be viewed as the associated tree of boundaries). It is not hard to see that any limit $\lim\Theta(G)$, viewed as a set, can be identied with appropriately chosen Martin’s set $\partial X$. (Namely, take as the underlying tree $T$ of $X$ the Bass-Serre tree $T_{G}$, take as the vertex spaces $X_{v}$ of $X$ the corresponding vertex groups $G(v)$, and as the edge spaces $X_{e}$ the corresponding edge stabilizers $G(e)$.) A less clear fact is that in case of $\lim\Theta(G)$ (viewed as special case of Martin’s $\partial X$) the topology on $\partial X$ described in Definitions 2.19 and 2.21 in [CM17] coincides with the topology on $\lim\Theta(G)$ provided in Section 2.3 of the present paper. We devote the remaining part of this proof to the justification of this coincidence of topologies, noting that then Theorem 3.2 follows fairly directly from Theorem 2.22 in [CM17].
Martin introduces his topology on $\partial X$ by means of describing families of neighbourhoods of points. (These neighbourhoods are not necessarily open in the resulting topology, but they contain open subneighbourhoods of the corresponding points.) In the case where $\partial X=\lim\Theta(G)$, we need to show that the identity map is a homeomorphism from $\partial X$ equipped with Martin’s topology to $\partial X$ equipped with our topology (as described in Section 2.3). However, since both these topological spaces are compact metrizable, it is sufficient to show that the identity map is continuous.
Suppose we are given a topological space $Z$ whose topology is given by a family $N(z):z\in Z$ of neighbourhood systems for points $z$, and a topological space $Z^{\prime}$ whose topology is induced by a basis $\mathcal{B}$ of open sets. Recall that then a map $f:Z\to Z^{\prime}$ is continuous iff for any open set $E\in{\mathcal{B}}$ and any $z\in f^{-1}(E)$ there is a neighbourhood $N\in N(z)$ of point $z$ such that $N\subset f^{-1}(E)$. In particular, if $Z=Z^{\prime}$ (as sets) then the identity map $\hbox{id}_{Z}:Z\to Z^{\prime}$ is continuous (with respect to the topologies mentioned above) iff for any $E\in{\mathcal{B}}$ and any $z\in E$ there is $N\in N(z)$ such that $N\subset E$.
Suppose now that $\partial X=\lim\Theta(G)$. Recall that $\lim\Theta(G)$, as a set, is the disjoint union of the following subsets: $\partial T_{G}$, $\Sigma_{u}:u\in V^{w}_{T_{G}}$ and $\dot{K}_{t}:t\in V^{b}_{T_{G}}$, where $\dot{K}_{t}=K_{t}\setminus\bigcup\{\Sigma_{u}:u\in N_{t}\}$. By a direct application of Definitions 2.19 and 2.21 of [CM17] in our context we get that Martin’s topology on $\partial X=\lim\Theta(G)$ is described by the following families of neighbourhoods of points:
  1. (1)
    for a point $\eta\in\partial T_{G}\subset\lim\Theta(G)$, and for any neighbourhood $U$ of $\eta$ in $V_{T_{G}}\cup\partial T_{G}$, we have a neighbourhood $V_{U}(\eta)$ of $\eta$ given by
    $V_{U}(\eta)=(U\cap\partial T_{G})\cup\bigcup\{\dot{K}_{t}:t\in U\}\cup\bigcup\{\Sigma_{u}:\{u\}\cup N_{u}\subset U\};$
  2. (2)
    for a point $\xi\in\dot{K}_{t}\subset\lim\Theta(G)$, and for any neighbourhood $U_{t}$ of $\xi$ in $K_{t}$, define $\hbox{Cone}_{\{U_{t}\}}(\xi)$ as the set consisting of all those $x\in V_{T_{G}}\cup\partial T_{G}\setminus\{t\}$ for which we have $\varphi_{e}(\Sigma_{w(e)})\subset U_{t}$, where $e$ is the last edge on the path from $x$ to $t$ in $T_{G}$; we then have a neighbourhood $V_{\{U_{t}\}}(\xi)$ of $\xi$ given by
    $V_{\{U_{t}\}}(\xi)=(\partial T_{G}\cap\hbox{Cone}_{\{U_{t}\}}(\xi))\cup\bigcup\{K_{s}:s\in\hbox{Cone}_{\{U_{t}\}}(\xi)\}\cup U_{t};$
  3. (3)
    for a point $\xi\in\Sigma_{u_{0}}\subset\lim\Theta(G)$, and for any family ${\mathcal{U}}=\{U_{u_{0}}\}\cup\{U_{t}:t\in N_{u_{0}}\}$ of neighbourhoods of $\xi$ in the spaces $\Sigma_{u_{0}}$ and $\{K_{t}:t\in N_{u_{0}}\}$, respectively, define $\hbox{Cone}_{\mathcal{U}}(\xi)$ as the set consisting of all those $x\in V_{T_{G}}\cup\partial T_{G}\setminus(\{u_{0}\}\cup N_{u_{0}})$ for which we have $\varphi_{e}(\Sigma_{w(e)})\subset U_{b(e)}$, where $e$ is the last edge on the shortest path in $T_{G}$ from $x$ to $N_{u_{0}}$; we then have a neighbourhood $V_{\mathcal{U}}(\xi)$ of $\xi$ given by
    $V_{\mathcal{U}}(\xi)=(\partial T_{G}\cap\hbox{Cone}_{\mathcal{U}}(\xi))\cup\bigcup\{K_{s}:s\in\hbox{Cone}_{\mathcal{U}}(\xi)\}\cup\big{[}\bigcup\{U_{t}:t\in N_{u_{0}}\}\setminus U_{u_{0}}\big{]}\cup$
    $\cup\big{[}U_{u_{0}}\cap\bigcap\{U_{t}:t\in N_{u_{0}}\}\big{]}.$
We are now ready for showing that the identity map of the set $\partial X=\lim\Theta(G)$, for the Martin’s topology in the source, and for the topology as in Section 2.3 in the target, is continuous. To do this, consider any open set $E\in{\mathcal{B}}$ (where $\mathcal{B}$ is as in Section 2.3), i.e. a set of form $E={\mathcal{G}}(W)$ for an appropriate $\Omega_{F}$-saturated open subset $W$ in a partial union $K_{F}$. We will proceed with the argument in the following three cases. First, let $\eta\in E$ be a point which belongs to $\partial T_{G}$. We need to find a neighbourhood of the form $N=V_{U}(\eta)$ as in (1) above, such that $N\subset E$. Let $t_{0}$ be this vertex of the b-subtree $F$ which is “closest” to $\eta$ in $T_{G}$, and let $e=[u_{0},t_{0}]$ be the last edge on the path in $T_{G}$ from $\eta$ to $t_{0}$. Put $U$ to be the set of all those elements $z$ in $(V_{T_{G}}\cup\partial T_{G})\setminus V_{F}$ for which the path in $T_{G}$ from $z$ to $t_{0}$ passes through $u_{0}$ (including $z=u_{0}$). $U$ is obviously a neighbourhood of $\eta$ in $V_{T_{G}}\cup\partial T_{G}$, and it is not hard to observe that the corresponding neighbourhood $V_{U}(\eta)$ is a subset of $E$, as required.
In the second case, let $\xi\in E$ be a point which belongs to $\dot{K}_{t}$, for some $t\in V^{b}_{T_{G}}$. We need to find a neighbourhood of the form $N=V_{\{U_{t}\}}(\xi)$ as in (2) above, such that $N\subset E$. Consider first the subcase when $t\notin V_{F}$, and let $e$ be the first edge on the path in $T_{G}$ connecting $t$ with the nearest vertex of $F$. Put $U_{t}=K_{t}\setminus\varphi_{e}(\Sigma_{w(e)})$. Obviously, $U_{t}$ is then a neighbourhood of $\xi$ in $K_{t}$, and the corresponding neighbourhood $V_{\{U_{t}\}}(\xi)$ is a subset of $E$, as required. In the second subcase, let $t\in V_{F}$. Observe that then $\xi\in W$. Let $e_{1},\dots,e_{k}$ be the edges of $F$ adjacent to $t$, and put $U_{t}=(K_{t}\cap W)\setminus\cup_{i=1}^{k}\varphi_{e_{i}}(\Sigma_{w(e_{i})})$. Again, it is easily seen that $U_{t}$ is a neighbourhood of $\xi$ in $K_{t}$, and that the corresponding neighbourhood $V_{\{U_{t}\}}(\xi)$ is a subset of $E$, as required.
In the final third case, let $\xi\in E$ be a point which belongs to $\Sigma_{u_{0}}$, for some $u_{0}\in V^{w}_{T_{G}}$. We need to find a neighbourhood of the form $N=V_{\mathcal{U}}(\xi)$ as in (3) above, such that $N\subset E$. We will do this by considering three subcases related to the possible position of $u_{0}$ with respect to $F$. First, consider the subcase where $u_{0}\in V_{F}$. Observe that then $\xi\in W$. Note also that, since $F$ is a b-subtree of $T_{G}$, the set $N_{u_{0}}$ of all neighbours of $u_{0}$ in $T_{G}$ is actually contained in $V_{F}$. For each $t\in N_{u_{0}}$, let ${\mathcal{E}}_{t}$ be the set of edges of $F$ adjacent to $t$ and distinct from $[u_{0},t]$. For each $t\in N_{u_{0}}$ put $U_{t}=(W\cap K_{t})\setminus\bigcup\{\varphi_{e}(\Sigma_{w(e)}):e\in{\mathcal{E}}_{t}\}$; put also $U_{u_{0}}=W\cap\Sigma_{u_{0}}$ and ${\mathcal{U}}=\{U_{u_{0}}\}\cup\{U_{t}:t\in N_{u_{0}}\}$. Note that for each $t\in N_{u_{0}}$, $U_{t}$ is a neighbourhood of $\xi$ in $K_{t}$, and similarly $U_{u_{0}}$ is a neighbourhood of $\xi$ in $\Sigma_{u_{0}}$. Moreover, it is not hard to see that the corresponding neigbourhood $V_{\mathcal{U}}(\xi)$ is then contained in $E$, as required. In the second subcase, we suppose that $u_{0}\notin V_{F}$, but $N_{u_{0}}\cap F\neq\emptyset$. For each $t\in N_{u_{0}}\setminus V_{F}$ put $U_{t}=K_{t}$. For $t\in N_{u_{0}}\cap V_{F}$, let ${\mathcal{E}}_{t}^{F}$ be the set of all edges in $F$ adjacent to $t$, and put $U_{t}=(W\cap K_{t})\setminus\bigcup\{\varphi_{e}(\Sigma_{w(e)}):e\in{\mathcal{E}}_{t}\}$. Put also $U_{u_{0}}=\Sigma_{u_{0}}$ and ${\mathcal{U}}=\{U_{u_{0}}\}\cup\{U_{t}:t\in N_{u_{0}}\}$. As before, note that for each $t\in N_{u_{0}}$, $U_{t}$ is a neighbourhood of $\xi$ in $K_{t}$, and similarly $U_{u_{0}}$ is a neighbourhood of $\xi$ in $\Sigma_{u_{0}}$. Likewise, it is not hard to see that the corresponding neigbourhood $V_{\mathcal{U}}(\xi)$ is then contained in $E$, as required. Finally, in the third subcase, we suppose that $N_{u_{0}}\cap F=\emptyset$. Denote by $t_{1}\in N_{u_{0}}$ this vertex which lies on the path from $u_{0}$ to $F$. Denote also by $u_{1}$ the white vertex adjacent to $t_{1}$ and distinct from $u_{0}$ which also lies on the path from $u_{0}$ to $F$. For any $t\in N_{u_{0}}\setminus\{t_{1}\}$, put $U_{t}=K_{t}$, and put $U_{t_{1}}=K_{t_{1}}\setminus\varphi_{[u_{1},t_{1}]}(\Sigma_{u_{1}})$. Put $U_{u_{0}}=\Sigma_{u_{0}}$ and ${\mathcal{U}}=\{U_{u_{0}}\}\cup\{U_{t}:t\in N_{u_{0}}\}$, and note again that $V_{\mathcal{U}}(\xi)$ is contained in $E$, as required.
This finishes the proof.
@@ 3.2. The reduced Bowditch JSJ splitting and its rigid cluster factors
Let $G$ be a 1-ended hyperbolic group which is not cocompact Fuchsian and let ${T}_{G}$ be the Bowditch JSJ tree of $G$. Let $\sim$ be the smallest equivalence relation on the vertex set of ${T}_{G}$ for which $u\sim v$ whenever $u$ and $v$ are rigid vertices at distance $2$. Equivalence classes of $\sim$ are either white vertex singletons, flexible vertex singletons or rigid classes (i.e. sets of rigid vertices). The convex hulls of the rigid classes are the rigid clusters of ${T}_{G}$. Each rigid cluster of ${T}_{G}$ is a subtree of $T_{G}$ (possibly a single vertex).
@@
We now describe closer the nature of the just defined reduced Bowditch JSJ splitting, and in particular of its factors. Observe that the bipartite structure of the tree ${T}_{G}$ descends to ${T}^{r}_{G}$, so that we can speak of black and white vertices of the latter. The white vertices of ${T}^{r}_{G}$ are naturally in bijective correspondence with those white vertices of ${T}_{G}$ which are adjacent to at least one flexible vertex. Their stabilizers (for the induced action of $G$ on ${T}^{r}_{G}$) coincide with the corresponding stabilizers for the $G$-action on ${T}_{G}$, and so they are maximal 2-ended subgroups of $G$. There are two kinds of black vertices in ${T}^{r}_{G}$. The first kind of black vertices of ${T}^{r}_{G}$ are those in bijective correspondence with the flexible vertices of ${T}_{G}$, and so we still call them flexible. Their stabilizers under the action of $G$ on ${T}^{r}_{G}$ do not change, and we call them the flexible factors of the reduced Bowditch JSJ splitting of $G$. The second kind of black vertices of ${T}^{r}_{G}$ are those in bijective correspondence with the rigid clusters of ${T}_{G}$, and we call them the rigid cluster vertices of ${T}^{r}_{G}$. Their stabilizers, which coincide with the stabilizers of the corersponding rigid clusters in the $G$-action on ${T}_{G}$, are called the rigid cluster factors of the reduced Bowditch JSJ splitting of $G$.
The next observations follow directly from the above description.
@@
Each rigid cluster vertex $v$ of the reduced Bowditch tree ${T}^{r}_{G}$ is isolated, i.e., there is no rigid cluster vertex in ${T}^{r}_{G}$ at distance 2 from $v$. Equivalently, all vertices of ${T}^{r}_{G}$ lying at distance 2 from a rigid cluster vertex $v$ are flexible.
@@
Given a rigid cluster $S\subset{T}_{G}$, and the corresponding rigid cluster factor $G(v_{S})$, consider the action of $G(v_{S})$ on $S$. The vertex stabilizers of this action are easily seen to coincide with the corresponding vertex stabilizers of the $G$-action on ${T}_{G}$. Thus, this action describes some splitting of the rigid cluster factor $G(v_{S})$, along maximal 2-ended subgroups, whose factors are some appropriate rigid factors of the Bowditch JSJ splitting of $G$. This observation is helpful for understanding the nature of rigid cluster factors, even though we will make no essential use of it.
Since the factors of any splitting of a hyperbolic group along quasi-convex subgroups are also quasi-convex (see e.g. Proposition 1.2 in [Bow98]), and since 2-ended subgroups are quasi-convex, we get the following.
@@
where each $\varphi_{e}:\partial G({w(e)})\to\partial G({b(e)})$ is the natural inclusion. The same argument as in the proof of Theorem 3.2 yields also the following result.
@@
Let $G$ be a 1-ended hyperbolic group which is not cocompact Fuchsian, and let $\Theta^{r}(G)$ be the associated tree system (related to the reduced Bowditch JSJ splitting of $G$). Then the Gromov boundary $\partial G$ is naturally homeomorphic to the limit of the tree system $\Theta^{r}(G)$, i.e.,
@@ 4. General trees of graphs
In this section we describe trees of graphs—a class of topological spaces which is the main object of study in this paper. As we will see, these spaces are compact metrizable, and the most interesting subclass (that of trees of 2-connected graphs) consists of spaces that are connected, locally connected, cutpoint-free, and of topological dimension 1. Throughout all of this section we make use of the terminology and notation introduced in Section 2. We start with basic definitions.
@@
for a tree system of punctured graphs (where for each $t\in V^{b}_{T}$, $\Gamma_{t}$ is some finite essential graph, i.e., nonempty and without isolated vertices).
@@ 4.1. Trees of graphs as inverse limits of graphs
We will now present an alternative description of trees of graphs, as inverse limits of some special inverse systems of graphs. (As we will see below, these inverse systems in fact coincide, up to isomorphism, with the systems ${\mathcal{S}}_{\Theta}$ as described in Section 2.6, associated to the corresponding tree systems $\Theta$ of punctured graphs.) This alternative approach to trees of graphs is related to (and generalizes) that given in Section 2 of [Ś21], where only a special case of the so called reflection trees of graphs has been treated, using a similar description.
Let $\Theta$ be a tree system of punctured graphs. For any graph $\Gamma_{t}$ whose punctured version $\Gamma_{t}^{\circ}$ appears as a constituent space in $\Theta$, denote by $D_{t}\subset|\Gamma_{t}|$ a countable dense subset of the realization $|\Gamma_{t}|$ containing all vertices of $\Gamma_{t}$. Note that such $D_{t}$ is unique up to a homeomorphism of $|\Gamma_{t}|$ that is the identity on the vertices and that preserves all the edges. Consider also the inverse system ${\mathcal{S}}_{\Theta}=[\{K_{F}^{\ast}\},\{f_{F_{1}F_{2}}\}]$ associated to $\Theta$. Observe that, viewing any black vertex $t\in V^{b}_{T}$ as a b-subtree of $T$, and denoting by $\{[\varphi_{[u,t]}(\Sigma_{u})]:u\in N_{t}\}$ the set of points of the quotient space $K_{t}^{\ast}$ obtained from $\Gamma_{t}^{\circ}$ by shrinking the peripheral sets $\varphi_{[u,t]}(\Sigma_{u}):u\in N_{t}$, there is a homeomorphism $h_{t}:K_{t}^{\ast}\to|\Gamma_{t}|$ such that:
@@
    $h_{t}$ respects types of the shrunk peripheral sets $[\varphi_{[u,t]}(\Sigma_{u})]$, i.e., if $\varphi_{[u,t]}(\Sigma_{u})$ is the boundary of a normal neighbourhood of some vertex of $\Gamma$, then $h_{t}([\varphi_{[u,t]}(\Sigma_{u})])$ is equal to this vertex, and if $\varphi_{[u,t]}(\Sigma_{u})$ is the boundary of a normal neighbourhood of some interior point of some edge $a$ of $\Gamma$, then $h_{t}([\varphi_{[u,t]}(\Sigma_{u})])$ is equal to some interior point of the same edge $a$.
@@
Let $X$ be the realization of a finite graph, and let $x\in X$. The blow-up of $X$ at $x$ is the space obtained from $X$ by deleting $x$, and by taking the completion of the resulting space which consists of adding as many points as the number of connected components into which $x$ splits its normal neighbourhood $B(x)$ in $X$. We denote this blow-up by $X^{\#}(x)$, and the set of points added at the completion by $P_{x}$ (we use the term blow-up divisor at $x$ for this set $P_{x}$). Similarly, if $J$ is a finite subset of $X$, the blow-up of $X$ at $J$, denoted $X^{\#}(J)$, is the space obtained from $X$ by performing blow-ups at all points of $J$ (in arbitrary order). Moreover, for any two finite subsets $J_{1}\subset J_{2}$ of $X$, denote by $\rho^{X}_{J_{2}J_{1}}:X^{\#}(J_{2})\to X^{\#}(J_{1})$ the blow-down map which maps the sets $P_{x}:x\in J_{2}\setminus J_{1}$ to the corresponding points $x$, and which is the identity on the remaining part of $X^{\#}(J_{2})$.
Now, let $F$ be a finite b-subtree of the underlying tree $T$ of the tree system $\Theta$. Recalling that we have identified graphs $|\Gamma_{t}|$ with the corresponding spaces $K_{t}^{*}$, for each $t\in V^{b}_{F}$ consider the subset $J_{t}^{F}\subset|\Gamma_{t}|$ given by $J_{t}^{F}:=\{h_{t}([\varphi_{[t,u]}(\Sigma_{u})]):u\in N^{F}_{t}\}$. Define $X_{F}$ as quotient of the disjoint union
@@
by the equivalence relation $\sim$ determined by the following gluing maps. Recall that, by condition (2) of Definition 4.1, for any white vertex $u$ of $F$ there are exactly two black vertices in $F$ adjacent to $u$. Denoting these adjacent black vertices by $t$ and $s$, consider the blow-up divisors $P_{h_{t}([\varphi_{[t,u]}(\Sigma_{u})])}\subset|\Gamma|^{\#}(J_{t}^{F})$ and $P_{h_{s}([\varphi_{[s,u]}(\Sigma_{u})])}\subset|\Gamma|^{\#}(J_{s}^{F})$. Note that there is a natural bijection